OpenStudy (anonymous):

Please help: Medal will be given. The volume of gas varies directly with temperature and inversely with pressure. Volume is 100 m3 when temperature is 1508 degrees and pressure is 15 lb/cm2. What is the volume when the temperature is 2508 degrees and the pressure is 20 lb/cm2?

OpenStudy (anonymous):

\[\frac{p_1 \times v_1}{T_1}=\frac{p_2 \times v_2}{T_2}\]

OpenStudy (anonymous):

where p1 is the pressure and v1 is the volume at T1 =1508 degrees

OpenStudy (anonymous):

can you just tell me if my answer is right 147.73m^3

OpenStudy (anonymous):

do you agree with the formula? have you done such a thing?

OpenStudy (anonymous):

its not letting me see your formula for some reason.. but i know v=kt/p

OpenStudy (anonymous):

\[\frac{100 \times 15}{1508}=\frac{20 \times v_2}{2508}\]

OpenStudy (anonymous):

100=k*1508/15 k=1500/1508 v= (1500/1508)*2508/20 =147.73m^3

OpenStudy (anonymous):

(p1x v1)/T1=(p2xv2)/T2

OpenStudy (anonymous):

can you see this formula now?

OpenStudy (anonymous):

yes your answer is correct

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

it is similar to my short cut concluded formula based on the definition of your question...

OpenStudy (anonymous):

you went and found the constant of proportionality then repeated the steps, i just made the whole thing in one steps, so they are the same, all roads lead to Rome...

OpenStudy (anonymous):

oh I see what your saying! so your saying your step is just easier then what i did? and my answer and steps are correct though?

OpenStudy (anonymous):

exactly...

OpenStudy (anonymous):

well i did not mean it is easier as much as i meant they are different ways to solve the same problem...

OpenStudy (anonymous):

okay I totally see what you are saying!! and my answer is correct right?

OpenStudy (anonymous):

yes if and only if you did not make an error with the calculator...

OpenStudy (anonymous):

is that what you got as an answer?

OpenStudy (anonymous):

so now you are more confident to recheck your answer again using the calculator...

OpenStudy (anonymous):

yes i can see the answer is logical from the figures...

OpenStudy (anonymous):

okay thank you for your help!!!

OpenStudy (anonymous):

well i am getting from computer calculator 124.73